OpenStudy (anonymous):

choose the ordered pair (x,y) that is a solution of the equation. 9x-13y=94 answers; (-3,-12) (-4,-10) (-6,-13) (-1,-7)

OpenStudy (anonymous):

@texaschic101 @jojo4eva @rose21 @baby1234567 @RANE

OpenStudy (texaschic101):

well...lets sub (-3,-12) 9(-3) - 13(-12) = 94 -27 + 156 = 94 129 = 94 (nope...not this one) (-4,-10) 9(-4) - 13(-10) = 94 -36 + 130 = 94 94 = 94 (correct) This is it...

OpenStudy (anonymous):

thanks again (x

OpenStudy (texaschic101):

anytime :)